# A Robust Large-Period Discrete Time Crystal and its Signature in a Digital Quantum Computer

### v2

@@ Results
Finally, we note that in our numerical results above, we have fixed the system size at $N=16$ so that it is larger than but remains close to that accessible in our quantum processor implementation below. However, as demonstrated in the Supplemental Materials, qualitatively similar results are also obtained at larger $N=32$.
Realization of robust $4T$-DTC on a quantum processor
@@
Figure 4: Physical signature of $4T$-DTC behavior on the IBM quantum processor and its comparison with tMPS results at several system sizes (green). The parameters used are: ${JT=0.16\pi}$, $MT=0.98\pi$, $hT=0.9\pi$. For the details of the device ibmq_cairo and its error information, see Fig. S5 in the Supplemental Materials.
@@
Here, we proceed to realize our system and capture its DTC signatures on the IBMQ quantum processor. Naive implementation of the time dynamics of our model in a quantum circuit follows from a similar trotterization procedure as in our tMPS simulations, and more details are shown in the Supplemental Materials. For any quantum circuit implementation, the coupling between qubits needs to be implemented via basic quantum gates such as the controlled NOT (CNOT) and the single-qubit rotation gates; one significant advantage of a quantum circuit implementation over other quantum platforms, e.g., ultracold atoms, is that a time-dependent model such as our DTC model can be implemented without any additional difficulty, just by concatenating different Trotter steps at different stages. Instead of transpiling local couplings within each Trotter step, we adopt a more efficient approach by leveraging the variational circuit optimization technique, and the details of this approach are presented in the Supplemental Materials. This strategy of directly implementing the whole circuit requires fewer $CX$ gates and compresses the circuit depth, thereby suppressing the effect of gate error. For other simulations using the Trotterization approach [68, 69], the signature of the DTC gets poorer after every Floquet cycle under the evolving dynamics, as the circuit depth itself grows linearly. In our unique scheme of trained circuits with a fixed circuit depth, as demonstrated below, we successfully achieve robust results for our proposed $4T$-DTC model, and in particular, it exhibits remarkably strong resilience to device noise even after many Floquet cycles. In the Supplemental Materials, we have also numerically simulated the system by using the direct Trotterization approach and found that a slightly smaller noise level than that in the actual NISQ device is required to observe a clear 4T-DTC signature. This in turn shows that our variational circuit optimization approach remains to be the more feasible approach to implement large-period DTCs on a present quantum computer.
@@
In Fig. 4, we present the stroboscopic magnetization $\langle S_{z}\rangle$ (solid circles) obtained from the IBM quantum computer over time dynamics and compare these with the numerical results (unfilled circles, squares, and triangles) obtained by the tMPS method. Remarkably, thanks to our variational method, the quantum simulation over long periods of time ($20$ Floquet steps) is possible. Within such long-time dynamics, our numerical and quantum results demonstrate an excellent agreement, indicating that our IBM Q simulation gives a perfect characterization of our $4T$-DTC model. Here, we execute our IBM Q simulation on an 8-qubit case which enables the realization of highly-compressed trained circuits for overcoming device noise  [^1] Our IBM Q simulation is scalable. Detailed discussion is shown in the Supplemental Materials.. This is already a sufficiently long chain for demonstrating 4T-DTC since we find that finite-size effects are qualitatively insignificant, i.e., our tMPS results at different sizes all show qualitatively similar profiles with the IBM Q results, both deep in the DTC regime and near the boundaries with the thermal phase (see Fig. 4 and S4 in the Supplemental Materials).
Discussion
@@
The experimental realization of our $4T$-DTC system at larger system sizes is expected to be a significant future research direction, both in the area of quantum computing and in condensed matter platforms such as ultracold atoms [70, 71, 72, 73, 74, 75, 76, 77, 78, 43, 79]. On the one hand, that our $4T$-DTC phase exists within a spin-1/2 system makes it a realistic and appropriate phenomenon for benchmarking the performance of various existing noisy intermediate-scale quantum (NISQ) devices. On the other hand, the ability to achieve a large size $4T$-DTC may also open up opportunities to harness its technological application beyond observing its subharmonic signatures, e.g., as a quantum memory or a passive quantum error correcting device. Finally, a realistic generalization of our spin-1/2 system construction that supports DTCs beyond period-quadrupling makes for a good avenue for future theoretical and experimental studies that can uncover rich phenomenology lying in the intersection of Floquet, many-body or even non-Hermitian physics.
Methods
Details of Suzuki-Trotter decomposition and time evolution on IBM Q
@@ Methods
^$\hat{\mathcal{H}}\left(t\right)=\begin{cases}-h/2\sum_{i=1}^{N_{0}}\left(H_{i}^{xx}-H_{i}^{yy}\right)-JH_{i}^{zz}&\quad 0<t<\frac{T}{2},\\ M\sum_{i=1}^{N_{0}}\sigma_{2i}^{x}&\quad\frac{T}{2}<t<T.\end{cases}$ (6)
@@
$\left[\prod_{i=0}^{N/2}e^{+i\delta th/2\left(\sigma_{2i}^{x}\otimes\sigma_{2i+1}^{x}-\left(1+\cos{\omega n\delta t}\right)\sigma_{2i}^{y}\otimes\sigma_{2i+1}^{y}\right)}\right],$
where $\delta t$ is the discretized time step, which is set to be $0.01/T$ ($T=1$) in our numerics. Note that for the next-nearest neighbor coupling terms in the above expression, for the simulation on IBM Q, it only requires a pair of swap gates between the ancilla qubit ($i=4$) and the physical qubit ($i=6$) with system size $N=8$ [see Fig. S5(d) in the Supplementary Materials], while for the simulation with tMPS, since it is aligned to a linear spin chain, for a system size of $N$ (assume $N$ is always even in this case), it requires $(N/2-1)$ pairs of the swap gates [65, 66]. Finally, the second-half time evolution operator $\hat{U}_{T/2\rightarrow T}$ can be simply realized by the following $R_{x}$ rotations
@@
Variational quantum algorithm for the time evolution on IBM Q
Here, we provide the essential details of our variational quantum circuit recompilation for the time evolution operators used in the main text. The scheme of our variational algorithm to obtain the parametrized quantum circuit is depicted in Fig. S5(c) and (d) in the Supplementary Materials. The original time evolution circuit $\hat{\mathcal{U}}$ with Suzuki-Trotter gates is transformed into a Trotterized ansatz circuit $\hat{\mathcal{V}}$ via variational optimization [see Fig. S5(c) in the Supplementary Materials]. For each stroboscopic time step $t=kT$ ($k=1,2,\cdots,\mathcal{N}_{\rm tot}$), it is transformed into an ansatz $\hat{\mathcal{V}}(t=kT)$, which consists of an initial layer of $U_{3}$ gates followed by concatenated odd layers (green) and even layers (purple)[Fig. S5(c) in the Supplementary Materials], which has a Trotterized time evolution pattern. Here, we set all number of total layers to be equal to be three, i.e. it has in total three combined layers of odd and even layers.
@@
[…]
Measurement of observables on IBM Q
[…]
@@
Data availability
The data that supports the findings of this study are available from the corresponding author upon request.
Code availability
The code used for this paper is available upon reasonable requests from the corresponding authors.
@@
R. W. .B. proposed the initial idea, and developed the $4T$-DTC model. C. H. L. and R. W. .B. guided the overall research direction and supervised this work. T. C. performed the classical simulations using tMPS, and also developed the algorithms and implemented them on the IBM Q. R. S. developed part of the algorithms on IBM Q and proposed the idea of the circuit recompilation. Y. B. contributed to part of the discussions. All the authors contributed to the writing of the manuscript.
@@ Acknowledgments
We are grateful to Jiangbin Gong for fruitful discussions. T. C. thanks E. Miles Stoudenmire for fruitful discussion via ITensor discourse group (https://itensor.discourse.group/). T. C. and R. S. thank Truman Ng and Russell Yang for discussions on the quantum simulation implementation on IBM Quantum services. C. H. L. and T. C. acknowledges support by Singapore’s NRF Quantum engineering grant NRF2021-QEP2-02-P09 and Singapore’s MOE Tier-II grant (MOE-T2EP50222-0003). T. C. and B. Y. acknowledges the support from Singapore National Research Foundation (NRF) under NRF fellowship award NRF-NRFF12-2020-0005. R. W. B acknowledges the support provided by the Deanship of Research Oversight and Coordination (DROC) at King Fahd University of Petroleum & Minerals (KFUPM) through project No. EC221010. We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team. The MPS calculation in this work is performed using ITensor library [80]. The computational work for this article was partially performed on resources of the National Supercomputing Centre, Singapore (https://www.nscc.sg/), and on the National University of Singapore (NUS)’s high-performance computing facilities.
Appendix S1 Robustness of $\pi/2$ quasienergy spacing in the presence of disorder

### v3

@@ 2.2 Evidence of robust $4T$-DTC via tMPS
We note that in our numerical results above, we have fixed the system size at $N=16$ so that it is larger than but remains close to that accessible in our quantum processor implementation below. However, as demonstrated in Appendix. B, qualitatively similar results are also obtained at larger $N=32$.
2.3 Finite size scaling analysis of the spectral properties of the many-body Floquet eigenstates
In Sec. 2.1, we have analytically elucidated the origin of our 4T-DTC model in terms of the spectral structure of the many-body Floquet eigenstates, i.e., they are long-range correlated and exhibit $\pi/2$-spectral gap. These features are exact and clearly seen at ideal parameter values whereby all many-body Floquet eigenstates can be analytically shown to take the form of macroscopic $4$-component cat states. Away from these ideal values, we argued that such features due to a mechanism reminiscent of quantum error correction. To further verify this argument and therefore the validity of $4T$-DTC, we shall numerically perform a thorough finite size scaling analysis of the aforementioned spectral properties in the following.
To quantify the long-range correlated nature of the many-body Floquet eigenstates, we define an Ising-ZZ spin-glass (SG) order parameter as [75, 76]
$\chi_{zz}=\mathcal{N}_{\chi}\sum_{\begin{subarray}{c}i,i^{\prime}\\ \gamma,\gamma^{\prime}\end{subarray}}\sum_{n=1}^{D}\left|\langle\epsilon_{n}|\sigma_{i,\gamma}^{z}\sigma_{i^{\prime},\gamma^{\prime}}^{z}|\epsilon_{n}\rangle\right|^{2},$ (5)
where $\mathcal{N}_{\chi}=\frac{2(N-2)!}{N!\;D}$ is the normalization factor ($N=2N_{0}$ is the total system size (number of qubits)), $D$ is the dimension of the total system Hilbert space, $|\epsilon_{n}\rangle$ represents an eigenstate of the Floquet (one-period) time evolution operator governed by the Hamiltonian from Eq. (1), and $\sigma_{i,\gamma}^{z}$ stands for the Pauli-$z$ operator resided on ladder $\gamma=a,b$ at site $i$. There, the first summation covers all the indices but excludes the cases $i=i^{\prime}$ and $\gamma=\gamma^{\prime}$, i.e., all the distinctive sites within the total system.
The case $\chi_{zz}=1$ signifies perfect long-range correlation with respect to $ZZ$ Pauli operator, whereas the other extreme case of $\chi_{zz}=0$ implies its absence. Our calculation (as shown in Fig. 4(a) (red triangle) at generic parameters deep in the DTC region) reveals that $\chi_{zz}$ takes an intermediate value between 0 and $1$, and it monotonically increases with the system size, thereby demonstrating the presence of long-range correlation in the many-body Floquet eigenstates with respect to $ZZ$ Pauli operator in the thermodynamic limit.
Figure 4: Finite size scaling analysis as a validity check of $4T$-DTC. (a) The Ising-ZZZZ spin-glass (SG) order $\chi_{zz}$ and the $\pi/2$-shifted spectral function $s_{\pi/2}$ as a function of the total system size; (b) The full power spectrum for different sizes from $N=4$ to $12$. All results are obtained by exact diagonalization of the Floquet unitary operator for the time evolution of the total system up to $t=400T$. In all panels, other parameters used are: $JT=0.16\pi,MT=1.0\pi$, and $hT=0.99\pi$.
To quantify the tendency of the many-body Floquet eigenstates to form $\pi/2$-spectral gap, we define another useful diagnostic termed the $\pi/2$-shifted spectral function $s_{\pi/2}$, which extends the $\pi$-shifted spectral function of Refs. [75, 76] for studying period-doubling DTCs. Specifically,
$s_{\pi/2}=\mathcal{N}_{s}\sum_{x\in\mathcal{M}}\int_{-\delta}^{\delta}\mathfrak{S}_{\hat{O},\pi/2}(\epsilon_{n},\eta)\,d\eta,$ (6)
where $\mathcal{N}_{s}=1/\sum_{n\in\mathcal{M}}\int_{-\pi/T}^{\pi/T}\mathfrak{S}_{\hat{O},\pi/2}(\epsilon_{n},\eta)\,d\eta$ is the normalization factor, $\mathcal{M}$ is a set of some random integers smaller than the the total system’s Hilbert space dimension  [^2] Throughout this work, we have chosen $\mathcal{M}=32$ for $N>4$, and $\mathcal{M}=12$ for $N=4$. , the integrand is defined as
$\mathfrak{S}_{\hat{O},\pi/2}=\sum_{k=-\infty}^{+\infty}\sum_{\epsilon_{m}}|\langle\epsilon_{n}|\hat{O}|\epsilon_{m}\rangle|^{2}$ (7)
$\times\delta\left[\epsilon_{n}-\epsilon_{m}-\eta-\left(2k+1/2\right)\pi/T\right],$
and $\hat{O}$ is an appropriately chosen operator that is expected to map a many-body Floquet eigenstate to another that is separated in quasienergy by $\pi/(2T)$. In our calculation, we have chosen $\hat{O}=\sigma_{a,0}^{z}+i\sigma_{b,0}^{z}$ as it yields the exact desired mapping at ideal parameter values. Indeed, at ideal parameter values ($J=0$ and $hT=\pi$), $s_{\pi/2}$ can be analytically computed to yield precisely $1$, which indicates that all many-body Floquet eigenstates exhibit perfect $\pi/2$ spectral gap. At more general parameter values, our numerical calculation of Fig. 4(a) (blue circles) reveals that $s_{\pi/2}$ takes a value between 0 and $1$, and it increases monotonically with the system size. This suggests that the many-body Floquet eigenstates indeed have the tendency to exhibit $\pi/2$-shifted spectral gap in the thermodynamic limit, and our operator choice $\hat{O}$ is relatively effective at capturing this spectral gap.
Finally, we have also performed a thorough analysis of the full power spectrum, calculated over a considerably long evolution time of up to $400T$, for various system sizes in Fig. 4(b). As expected, the subharmonic peaks at $1/4$ and $3/4$ the driving frequency increase monotonically with the system size, thereby highlighting the lifetime improvement of the period-quadrupling behavior with the system size.
To conclude, our numerics have confirmed that our system represents a genuine $4T$-DTC based on its three defining characteristics that persist at general parameter values: (i) the many-body Floquet eigenstates are long-range correlated ($\chi_{zz}$ is finite and increases monotonically with the system size), (ii) the corresponding Floquet quasienergies exhibit $\pi/2$-shifted spectral gap ($s_{\pi/2}$ is finite and increases monotonically with the system size), and (iii) the period-quadrupling behavior lasts longer with increase in the system size (the subharmonic peaks at $1/4$ and $3/4$ the driving frequency increase monotonically with the system size).
2.4 Realization of robust $4T$-DTC on a quantum processor
@@
Figure 5: The stroboscopic magnetization ($\langle S_{z}\rangle$) obtained from an IBM quantum processor (solid blue circles) and tMPS at several system sizes (empty square/triangular marks) as a function of the number of periods ($t/T$). The parameters used are: $JT=0.16\pi$}, $MT=0.98\pi$, $hT=0.9\pi$. For the details of the device ibmq_cairo and its error information, see Fig. 11 in Appendix. D.4.
@@
Here, we proceed to realize our system and capture its DTC signatures on the IBM quantum processor. Naive implementation of the time dynamics of our model in a quantum circuit follows from a similar trotterization procedure as in our tMPS simulations, and more details are shown in the Methods and Appendix. For any quantum circuit implementation, the coupling between qubits needs to be implemented via basic quantum gates such as the controlled-NOT (CNOT) and the single-qubit rotation gates; one significant advantage of a quantum circuit implementation over other quantum platforms, e.g., ultracold atoms, is that a time-dependent model such as our DTC model can be implemented without any additional difficulty, just by concatenating different Trotter steps at different stages. Instead of transpiling local couplings within each Trotter step, we adopt a more efficient approach by leveraging the variational circuit optimization technique, and the details of this approach are presented in Appendix. D. This strategy of directly implementing the whole circuit requires fewer $CX$ gates and compresses the circuit depth, thereby suppressing the effect of gate error. For other simulations using the Trotterization approach [77, 78], the signature of the DTC gets poorer after every Floquet cycle under the evolving dynamics, as the circuit depth itself grows linearly. In our unique scheme of trained circuits with a fixed circuit depth, as demonstrated below, we successfully achieve robust results for our proposed $4T$-DTC model, and in particular, it exhibits remarkably strong resilience to device noise even after many Floquet cycles. In Appendix. D.4, we have also numerically simulated the system by using the direct Trotterization approach and found that a slightly smaller noise level than that in the actual NISQ device is required to observe a clear 4T-DTC signature. This in turn shows that our variational circuit optimization approach remains to be the more feasible approach to implement large-period DTCs on a present quantum computer.
@@
In Fig. 5, we present the stroboscopic magnetization $\langle S_{z}\rangle$ (solid circles) obtained from the IBM quantum computer over time dynamics and compare these with the numerical results (unfilled circles, squares, and triangles) obtained by the tMPS method. Remarkably, thanks to our variational method, the quantum simulation over long periods of time ($20$ Floquet steps) is possible. Within such long-time dynamics, our numerical and quantum results demonstrate an excellent agreement, indicating that our IBM Q simulation gives a perfect characterization of our $4T$-DTC model. Here, we execute our IBM Q simulation on an 8-qubit case which enables the realization of highly-compressed trained circuits for overcoming device noise  [^3] Our IBM Q simulation is scalable. Detailed discussion is shown in tAppendix. D.1.. This is already a sufficiently long chain for demonstrating 4T-DTC since we find that finite-size effects are qualitatively insignificant, i.e., our tMPS results at different sizes all show qualitatively similar profiles with the IBM Q results, both deep in the DTC regime and near the boundaries with the thermal phase (see Fig. 5 and 9 in Appendix. C). While a strict phase of matter is formally defined in the thermodynamic limit, it is standard practice to infer phase behavior from finite-sized systems when its key signatures persist over several different system sizes. Indeed, our results, supported by both simulations (Sec. 2.3) and experiments (Sec. 2.4), clearly demonstrate robust subharmonic response and other expected characteristics of the 4T-DTC phase at varying system sizes, indicating that the observed phenomena are not finite-size artifacts but actual features of a genuine non-equilibrium phase.
3 Discussion
@@
Besides in the area of condensed matter systems, the physical demonstration of our $4T$-DTC system at larger system sizes on a digital quantum computer is expected to be a significant future research direction in the area of quantum computing, particularly on the topics of time crystals [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. On the one hand, that our $4T$-DTC phase exists within a spin-1/2 system makes it a realistic and appropriate phenomenon for benchmarking the performance of various existing noisy intermediate-scale quantum (NISQ) devices. On the other hand, the ability to achieve a large size $4T$-DTC may also open up opportunities to harness its technological application beyond observing its subharmonic signatures, e.g., as a quantum memory or a passive quantum error correcting device. Finally, a realistic generalization of our spin-1/2 system construction that supports DTCs beyond period-quadrupling makes for a good avenue for future theoretical and experimental studies that can uncover rich phenomenology lying in the intersection of Floquet, many-body or even non-Hermitian physics such as in ultracold atomic systems [79, 80, 81, 82, 83, 84, 85, 86, 87, 47, 88].
4 Methods
4.1 Details of Suzuki-Trotter decomposition and time evolution on IBM Q
@@
$\hat{\mathcal{H}}\left(t\right)=\begin{cases}-h/2\sum_{i=1}^{N_{0}}\left(H_{i}^{xx}-H_{i}^{yy}\right)-JH_{i}^{zz}&\quad 0<t<\frac{T}{2},\\ M\sum_{i=1}^{N_{0}}\sigma_{2i}^{x}&\quad\frac{T}{2}<t<T.\end{cases}$ (9)
@@
[…]
where $\delta t$ is the discretized time step, which is set to be $0.01/T$ ($T=1$) in our numerics. Note that for the next-nearest neighbor coupling terms in the above expression, for the simulation on IBM Q, it only requires a pair of swap gates between the ancilla qubit ($i=4$) and the physical qubit ($i=6$) with system size $N=8$ [see Fig. 11 in Appendix. D.4], while for the simulation with tMPS, since it is aligned to a linear spin chain, for a system size of $N$ (assume $N$ is always even in this case), it requires (N/21)(N/2-1) pairs of the swap gates [72, 73]. Finally, the second-half time evolution operator $\hat{U}_{T/2\rightarrow T}$ can be simply realized by the following $R_{x}$ rotations
@@
4.2 Variational quantum algorithm for the time evolution on IBM Q
Here, we provide the essential details of our variational quantum circuit recompilation for the time evolution operators used in the main text. The scheme of our variational algorithm to obtain the parametrized quantum circuit is depicted in Fig. 11(c) and (d) in Appendix. D.4. The original time evolution circuit $\hat{\mathcal{U}}$ with Suzuki-Trotter gates is transformed into a Trotterized ansatz circuit $\hat{\mathcal{V}}$ via variational optimization [see Fig. 11(c) in Appendix. D.4]. For each stroboscopic time step $t=kT$ ($k=1,2,\cdots,\mathcal{N}_{\rm tot}$), it is transformed into an ansatz $\hat{\mathcal{V}}(t=kT)$, which consists of an initial layer of $U_{3}$ gates followed by concatenated odd layers (green) and even layers (purple) [Fig. 11(c) in Appendix. D.4], which has a Trotterized time evolution pattern. Here, we set all number of total layers to be equal to be three, i.e. it has in total three combined layers of odd and even layers.
@@ 4.2 Variational quantum algorithm for the time evolution on IBM Q
where we have deliberately chosen the initial $|\psi_{0}\rangle$ with all sites in $|\uparrow\rangle$. The process of the circuit variational optimization involves optimizing a $3D$ rotation gate labeled as $U_{3}(\theta,\phi,\lambda)$. This gate is characterized by three rotational parameters which are variable: $\theta$, $\phi$ and $\lambda$. We refer to the Appendix. D.2 for other details of the optimization.
4.3 Measurement of observables on IBM Q
The magnetization in the $z$ direction for each spin residing on ladder a from our model is obtained via the measurement procedure on IBM quantum processor, which is performed after the time evolution in the Trotterized ansatz circuit $\hat{\mathcal{V}}$. On IBM Q, the measured outcomes are all represented in binary bit strings, i.e. 0 for spin-up ($|\uparrow\rangle$), and $1$ for spin-down ($|\downarrow\rangle$). For each site $i$, the magnetization in the $z$ direction for each spin residing on ladder a $\langle\sigma_{i,a}^{z}\rangle$ is computed as
@@ 4.3 Measurement of observables on IBM Q
Code and data availability
We have uploaded the essential codes and data which are related to this work at [89].
@@
R. W. B. proposed the initial idea, and developed the $4T$-DTC model. C. H. L. and R. W. .B. guided the overall research direction and supervised this work. T. C. performed the classical simulations using tMPS, and also developed the algorithms and implemented them on the IBM Q. R. S. developed part of the algorithms on IBM Q and proposed the idea of the circuit recompilation. Y. B. contributed to part of the discussions. All the authors contributed to the writing of the manuscript.
@@ Acknowledgments
We are grateful to Jiangbin Gong and Tim Byrnes for fruitful discussions. T. C. thanks E. Miles Stoudenmire for fruitful discussion about ITensor usage via ITensor discourse group (https://itensor.discourse.group/). T. C. and R. S. thank Truman Ng and Russell Yang for discussions on the quantum simulation implementation on IBM Quantum services. C. H. L. and T. C. acknowledges support by Singapore’s NRF Quantum engineering grant NRF2021-QEP2-02-P09 and Singapore’s MOE Tier-II grant (MOE-T2EP50222-0003). T. C. and B. Y. acknowledges the support from Singapore National Research Foundation (NRF) under NRF fellowship award NRF-NRFF12-2020-0005. R. W. B acknowledges the support provided by the Deanship of Research Oversight and Coordination (DROC) at King Fahd University of Petroleum & Minerals (KFUPM) through project No. EC221010. We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team. The MPS calculation in this work is performed using ITensor library [90]. The computational work for this article was partially performed on resources of the National Supercomputing Centre, Singapore (https://www.nscc.sg/). We are also gratefully acknowledge computational resources from the A*STAR Computational Resource Center (ACRC). T. C. also acknowledge that this work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise, made possible through the collaboration and support of Tim Byrnes.
Appendix A Robustness of $\pi/2$ quasienergy spacing in the presence of disorder